OpenStudy (anonymous):

HELP HELP HEEELP? :{ The lengths of two sides of a right triangle are given. Find the length of the third side. Round to the nearest tenth if necessary. legs: 18 in. and 17 in. a. 24.8 in. b. 5.9 in. c. 27.7 in. d. 19.5 in.

OpenStudy (mathstudent55):

Do you know the Pythagorean theorem?

OpenStudy (mathstudent55):



OpenStudy (anonymous):

a^2 + b^2 = c^2 so 18^2+17^2=c^2 find what 18^2 plus 17^2 is and the take the square root of the number given and theres your answer

OpenStudy (mathstudent55):

You are given a and b, and you need to solve for c.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

thanks. i forgot how to do these.

OpenStudy (mathstudent55):

You're welcome.